The G.F packs 'shrimphs' in three categories:"small", Medium and Large, with average weight of 1.4kgs and standard deviation of 0.6kgs. if a pack weight more than 1.91kgs, it is marked 'large'. if a pack weight less than 0.47kgs, it is marked as small.A pack weighing in between 0.47kgs to 1.91kgs is marked as medium. identify the proportion of packs in each category:
weight: more than 1.91kgs between 1.91kgs & .47kgs less than .47kgs
category: large medium small
Prop. of Packs: - - -

Answer by stanbon(75887)

The G.F packs 'shrimp' in three categories:"small", Medium and Large, with average weight of 1.4kgs and standard deviation of 0.6kgs.
-----------------------
if a pack weighs more than 1.91kgs, it is marked 'large'.
if a pack weight less than 0.47kgs, it is marked as small
A pack weighing in between 0.47kgs to 1.91kgs is marked as medium.
---------------------
identify the proportion of packs in each category:
weight: more than 1.91kgs
between 1.91kgs & .47kgs
less than .47kgs
------
I'll work one part and leave the other two to you.
Proportion of medium:
Find z(0.47) = (0.47-1.4)/0.6 = -1.55
Find z(1.91) = (1.91-1.4)/0.6 = 0.85
----
Prop of medium = P(-1.55<= z <=0.85) = normalcdf(-1.55,0.85) = 0.7418
=================
Cheers,
Stan H.
category: large medium small
Prop. of Packs: - - -